12-67
mated residual value $10,000 1,000Depreciation m In 2020, Bellico increased the estimated residual value to $2,000 five years for financial accounting purposes. Additional information follows and increased the total Expenses other than depreciation and tax...... 25,000 Required a. Provide the 2019 entry for depreciation and the ending 2019 accumulated depreciation bolance. b. Provide the 2020 entry for depreciation and the ending 2020 accumulated depreciation balance. c. Provide the comparative 2019 and 2020 income statements, including disclosures related to the accounting estimate. Pety Corporation has been depreciating equipment over a 10-year life on a straighi-line basis The which costs $24,000, was purchased on January 1, 2016. The equipment has an estimated residual vale $6,000. On the basis of experience since acquisition, management has decided to depreciate the a total life of 14 years instead of 10, with no change in the estimated residual value. The change is to ef tive on January 1, 2020. The annual financial statements are prepared on a comparative basis (2019 and 3020 m presented). 2019 income and 2020 income before depreciation for 2019 and 2020 were $49,.800 and $52,80, respectively. Disregard income tax considerations. Exercise 12-67 Recording and On in Estimated Useful Life LO4 24, 000 Required a. Prepare the journal entries to be recorded in 2016-2019 for annual depreciation. Determine the ending bal- ance in accumulated depreciation on December 31, 2019. b. Prepare the entry to record depreciation expense in 2020, the year of the change. c. Illustrate how the change should be reported on the 2020 balance sheet and income statement, w tstanding are 100,00), -of-the-years-digis tha heninning of 2020, with no change in he companied by the 2019 results for comparative purposes (shares of common stock ou Southern Corporation has been depreciating equipment over a 10-year life using the sum method. The equipment was acquired January 1, 2016, and cost $68,000 (estimated residual value, $13,000).The xercise 12-68
Answer #1
a. Depreciation expense using straight line method=(Cost - Residual value)/Useful life=(24000-6000)/10=$ 1800
Journal entry
Year General journal Debit Credit
2016 Depreciation expense 1800
Accumulated depreciation 1800
2017 Depreciation expense 1800
Accumulated depreciation 1800
2018 Depreciation expense 1800
Accumulated depreciation 1800
2019 Depreciation expense 1800
Accumulated depreciation 1800
Accumulated depreciation on Dec 31,2019=1800*4=$ 7200
b. Depreciation expense for 2020 onwards=(Book value-residual value)/remaning life of asset
Book value=Cost-Accumulated depreciation=24000-7200=$ 16800
Remaining life of asset=Revised usegul life-life expired=14-4=10 years
Depreciation expense for 2020 onwards=(16800-6000)/10=$ 1080
Journal entry
Year General journal Debit Credit
2016 Depreciation expense 1080
Accumulated depreciation 1080
c. Income statement (Extract)
2020 2019
Income before depreciation 52800 49800
Less: depreciation 1800 1080
Net income 51000 48720
Earnings per share 0.51 0.49
(51000/100000) (48720/100000)
Balance sheet (Extract)
2020 2019
Property,plant and equipment:
Equipment
At cost 24000 24000
Less: Accumulated depreciation 8280 7200
(7200+1080)
Net value 15720 16800